For the vector space P2,find the dual basis of {1+x,1+2x,1+x+x2}?
1
Expert's answer
2020-02-25T08:37:11-0500

Dual basis:

f1(x)=a1+a2*x+a3*x^2

f2(x)=b1+b2*x+b3*x^2

f3(x)=c1+c2*x+c3*x^2


y1=1+x

y2=1+2x

y3=1+x+x^2


fi(yi)=1; fi(yj)=0 (i "\ne" j)


1 = a1+a2

0 = b1+b2

0 = c1+c2

1 = b1+2b2

0 = a1+2a2

0 = c1+2c2

1 = c1+c2+c3

0 = a1+a2+a3

0 = b1+b2+b3


Solving this system we obtain

a1=2, a2=-1, a3=-1, b1=-1, b2=1, b3=0, c1=0, c2=0, c3=1


Answer:

f1=2-x-x^2

f2=-1+x

f3=x^2